OpenStudy (anonymous):

Differentiate y= (2x-3)^4 (x^2+x+1)^5

OpenStudy (espex):

You need to use the chain and product rules, then simplify. \[y{\prime} =f(x)g{\prime}(x) + f{\prime}(x)g(x)\] Where \(g(x) = (x^2+x+1)^5\) and \(f(x) = (2x-3)^4\)